OpenStudy (rianiscool):

will give a medal this is math boo boo math !! 00 [][] Evaluate the expression. 7 · (8 + 3) A. 18 B. 29 C. 59 D. 77

OpenStudy (anonymous):

D

OpenStudy (rianiscool):

how

OpenStudy (rianiscool):

who ever tell me why it is that letter i will give the medal to

OpenStudy (igreen):

This is very easy. 7 · (8 + 3) 7 · (8 + 3) = 7(8 + 3) Distribute 7 into the parenthesis by multiplying 7 to every number in the parenthesis: \(7 * 8 = 56\) \(7 * 3 = 21\) So we have 56 + 21, which is 77.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

8+3=11 11*7=77

OpenStudy (rianiscool):

woooooooooowwwww igreen

OpenStudy (anonymous):

Remember PEMDAS!

OpenStudy (igreen):

^

OpenStudy (anonymous):

Because you do PEMDAS

OpenStudy (rianiscool):

thank you

OpenStudy (rianiscool):

pemdas

OpenStudy (rianiscool):

so d

OpenStudy (igreen):

PEMDAS P = Parenthesis first (unless there isn't a parenthesis) E = Exponent (you do this first if there's no parenthesis or if it's in the parenthesis) M & D = Multiplication and Division (you do whichever comes first from left to right) A & S = Addition and Subtraction (you do whichever comes first from left to right)

OpenStudy (igreen):

Yes, @rianiscool

OpenStudy (rianiscool):

you get the medal igreen